OpenStudy (anonymous):

Carlee wants to rent a car for a week and pay no more than $150. If the rental company charges $80 per week plus $.18 per mile. What is the greatest number of miles she can drive the rental car and stay within her budget?

OpenStudy (anonymous):

fixed cost 150 - 80= 70 70 / 0.20 = 350 miles

OpenStudy (anonymous):

388 miles

OpenStudy (anonymous):

yeah your right, i rounded it to 0.20 cents

OpenStudy (anonymous):

150 - 80 = 70 70/ .18 = 388.8 which equals to 388 miles the maximum

OpenStudy (anonymous):

thanks